OpenStudy (anonymous):

Simplify ^5sqrt(- frac{1}{3125})

OpenStudy (anonymous):

\[^5\sqrt (\frac{1}{3125} )\] Wasnt sure how to write this, but everything in the sqrt is in it.

OpenStudy (anonymous):

0.2

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

take the fifth root of the top and bottom. Realize that 3125=5^5 so the fifth root of 3125 is 5. Then you have 1/5